### v2

# Neural Architecture Search via Two Constant Shared Weights Initialisations

Abstract
In recent years, zero-cost metrics are gaining ground in neural architecture search (NAS). These metrics allow finding the optimal neural network for a given task faster and with a lesser computational load than conventional NAS methods. Equally important is that they also shed some light on the internal workings of neural architectures. This paper presents a zero-cost metric that highly correlates with the train set accuracy across the NAS-Bench-101, NAS-Bench-201 and NAS-Bench-NLP benchmark datasets. We evaluate a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. For this, we only use an unlabelled mini-batch of data. We observe that the dispersion of the outputs between two initialisations positively correlates with trained accuracy. The correlation further improves when we normalise dispersion by average output magnitude. The resulting metric, epsilon, does not require gradients computation and unbinds the NAS procedure from training hyperparameters, loss metrics and human-labelled data. Our method is easy to integrate within existing NAS algorithms and takes a fraction of a second to evaluate a single network. The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
Keywords: machine learning, neural architecture search, zero-cost NAS
1 Introduction
The field of neural architecture search (NAS) emerged about a decade ago as an effort to automatise the process of neural geometry optimisation. At the early stages of NAS development, every candidate architecture used to be evaluated through the training process (reinforcement learning (Williams (1992)), evolutionary algorithms (Real et al. (2019)), Bayesian optimisation (Falkner et al. (2018); White et al. (2021))).
One-shot algorithms adopting weight sharing dispense of multiple architectures training, reducing the search time drastically (efficient reinforcement learning (Pham et al. (2018)), random search with parameters sharing (Li and Talwalkar (2020)), differentiable methods). Nevertheless, they require the training of a massive hypernet, which necessitates elaborate hyperparameter tuning. While these methods prove efficient, they do not systematically achieve satisfactory results (Dong and Yang (2019)). One of the best of them, DARTS- (Chu et al. (2020)), shows significant uncertainty compared to evolutionary or reinforcement algorithms.
There are available methods that estimate network performance without training the dataset of interest but relying on an auxiliary predictive machine learning (ML) model built on a dataset of trained architectures (Istrate et al. (2019); Deng et al. (2017)). These methods aim to accelerate the NAS process for image recognition but still rely on training and cannot apply to other ML problems.
[…]
1.1 Zero-cost NAS
To alleviate the process of architecture search, many researchers focus on developing methods that allow finding optimal architectures without model training – so-called zero-cost NAS methods. These methods evaluate networks via a trainless proxy. They typically require the equivalent of one or a few training epochs, which is two to three orders of magnitude faster than previous NAS methods.
Weight agnostic neural networks. One of the pioneering works in zero-shot NAS is presented by Gaier and Ha (2019). The authors demonstrate a constructor building up neural architectures based on the mean accuracy over several initialisations with constant shared weights and the number of parameters contained within the model. The resulting model achieves over $90\%$ accuracy on MNIST data (LeCun et al. (2010)) when the weights are fixed to the best-performing constants. While these results are very intriguing, the authors admit that such architectures do not perform particularly well upon training. Moreover, back in $2019$, the benchmark databases of trained architectures, which are now routinely used to compare NAS metrics with each other, were yet to be released, which disables the comparison of this zero-shot method against the most recent ones.
ReLU activation pattern. In $2020$, Mellor et al. (2020) present the naswot metric, which exploits the rectified linear unit (ReLU, Agarap (2018)) activation function’s property to yield distinct activation patterns for different architectures. Concretely, every image yields a binary activation vector upon passing through a network, forming a binary matrix for a mini-batch. The logarithm of the determinant of this matrix serves as a scoring metric. Authors show that larger naswot values are associated with better training performances, which leads to the conclusion that high-performing networks should be able to distinguish the inputs before training. Unfortunately, the method can only be implemented on networks with ReLU activation functions, which limits its applicability to convolutional architectures.
Another work employing the aforesaid ReLU property is the one from Chen et al. (2021). They combine the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to build the tenas metric. Instead of evaluating each network in the search space individually, they create a super-network built with all the available edges and operators and then prune it.
Coefficient of variance. Our previous work on fully trainless NAS (Gracheva (2021)) evaluates the stability of untrained scores over random weights initialisations. The networks are initialised with multiple random seeds, and architectures are selected based on the coefficient of variance of the accuracy at initialisation, cv. While cv performance is associated with a high error rate, the author concludes that a good architecture should be stable against random weight fluctuations. While this method can, in theory, apply to any neural architecture type, it requires multiple initialisations and is relatively heavy compared to naswot and later methods. Furthermore, accuracy-based scoring metrics can only apply to classification problems, and it is unclear how to extend cv implementation to the regression tasks.
Gradient sign. The grad_sign metric is built to approximate the sample-wise optimisation landscape (Zhang and Jia (2021)). The authors argue that the closer local minima for various samples sit to each other, the higher the probability that the corresponding gradients will have the same sign. The number of samples yielding the same gradient sign approximates this probability. It allows to evaluate the smoothness of the optimisation landscape and architecture trainability. The method requires labels and gradient computation.
Pruning-at-initialisation proxies. Several powerful zero-cost proxies have emerged as an adaptation of pruning-at-initialisation methods to NAS in the work by Abdelfattah et al. (2021): grad_norm (Wang et al. (2020)), snip (Lee et al. (2018)), synflow (Tanaka et al. (2020)). These metrics are originally developed to evaluate the network’s parameters’ salience and prune away potentially meaningless synapses. They require a single forward-backwards pass to compute the loss. Then, the importance of parameters is computed as a multiplication of the weight value and gradient value. Then, they integrate the salience over all the parameters in the network to evaluate its potential upon training. What is particularly interesting about the synflow metric is that it evaluates the architectures without looking at the data by computing the loss as the product of all the weights’ values (randomly initialised). synflow metric shows the most consistent performance among various search spaces and sets the state-of-the-art for the zero-cost NAS.
Both naswot and synflow do not depend on labels, which arguably reduces the effect of human error during data labelling. Moreover, naswot does not require gradient computation, which renders this method less memory-intensive.
The above results imply that neural networks might have some intrinsic property which defines their prediction potential before training. Such property should not depend on the values of trainable parameters (weights) but only on the network’s topology. In the present work, we combine the takeaways from the existing trainless NAS implementations to present a new metric which significantly outperforms existing zero-cost NAS methods.
1.2 NAS benchmarks
To guarantee our metric’s reproducibility and compare its performance against other NAS algorithms, we evaluate it on the three widely used NAS benchmark datasets.
NAS-Bench-101. The first and one of the largest NAS benchmark sets of trained architectures. It consists of $423{,}624$ cell-based convolutional neural networks (Ying et al. (2019)). The architectures consist of three stacks of cells, each followed by max-pooling layers. Cells may have up to $7$ vertices and $9$ edges, with $3$ possible operations. This benchmark is trained multiple times on a single dataset, CIFAR-10 (Krizhevsky et al. (2009)), with a fixed set of hyperparameters for $108$ epochs.
NAS-Bench-201. It is a set of architectures with a fixed skeleton consisting of a convolution layer and three stacks of cells connected by a residual block (Dong and Yang (2020)). Each cell is a densely connected directed acyclic graph with $4$ nodes, $5$ possible operations and no limits on the number of edges, providing a total of $15{,}625$ possible architectures. Architectures are trained on three major datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al. (2009)) and a downsampled version of ImageNet (Chrabaszcz et al. (2017)). Training hyperparameters are fixed, and the training spans $200$ epochs.
NAS-Bench-NLP. As the name suggests, this benchmark consists of architectures suitable for neural language processing (Klyuchnikov et al. (2022)). Concretely, it consists of randomly-generated recurrent neural networks. Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations. Here, we only consider models trained and evaluated on Penn Treebank (PTB, Marcinkiewicz (1994)) dataset: $14{,}322$ random networks with a single seed and $4{,}114$ with tree seeds. The training spans $50$ epochs and is conducted with fixed hyperparameters.
2 Epsilon metric
Two existing NAS methods inspire the metric that we share in the present work: cv (Gracheva (2021)) and weight agnostic neural networks (Gaier and Ha (2019)). Both metrics aim to exclude individual weight values from consideration when evaluating networks: the former cancels out the individual weights via multiple random initialisations, while the latter sets them to the same value across the network. It is very intriguing to see that an network can be characterised purely by its topology.
As mentioned above, the cv metric has two principal disadvantages. While it shows a fairly consistent trend with trained accuracy, it suffers from high uncertainty. It can be, to some degree, explained by the fact that random weight initialisations bring in some noise. Our idea is that replacing random initialisations with single shared weight initialisations, similarly to Gaier and Ha (2019), should improve the method’s performance.
The second weak point is that cv is developed for classification problems and relies on accuracy. Therefore, it needs to be clarified how to apply this metric to regression problems. The coefficient of variation is a ratio of standard deviation to mean, and Gracheva (2021) shows that cv correlates negatively with train accuracy. It implies that the mean untrained accuracy should be maximised. On the other hand, for regression tasks, performance is typically computed as some error, which is sought to be minimised. It is not apparent whether the division by mean untrained error would result in the same trend for cv metric.
To address this issue, we decided to consider raw outputs. This modification renders the method applicable to any neural architecture. However, it comes with a difference: accuracy returns a single value per batch of data, while raw outputs are $[N_{\textrm{BS}}\times L]$ matrices, where $N_{\rm{BS}}$ is the batch size and $L$ is the length of a single output. [^1] This length depends on the task and architecture: for regression tasks $L=1$, for classification, $L$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string. In our work, we flatten these matrices to obtain a single vector $\mathbf{v}$ of length $L_{v}=N_{\textrm{BS}}\times L$ per initialisation. We then stuck both initialisations into a single output matrix $\mathbf{V}$.
Before proceeding to statistics computation over initialisations, we also must normalise the output vectors: in the case of constant shared weights, outputs scale with weight values. In order to compare initialisations on par with each other, we use min-max normalisation:
$\mathbf{V}_{i}^{\prime}=\frac{\mathbf{V}_{i}-\min(\mathbf{V}_{i})}{\max(\mathbf{V}_{i})-\min(\mathbf{V}_{i})},$ (1)
where $i$ is the index for initialisations, $i=\{0,1\}$.
We noticed that two distinct weights are sufficient to grasp the difference between initialisations. Accordingly, instead of standard deviation, we use mean absolute error between the normalised outputs of two initialisations:
$\delta=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\lvert\mathbf{V}_{1,j}^{\prime}-\mathbf{V}_{2,j}^{\prime}\rvert.$ (2)
The mean is computed over the outputs of both initialisations as follows:
$\mu=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\frac{\mathbf{V}_{1,j}^{\prime}+\mathbf{V}_{2,j}^{\prime}}{2}=\frac{1}{2L_{v}}\sum\limits_{i=0}^{2}\sum\limits_{j=0}^{L_{v}}\mathbf{V}_{i,j}^{\prime}$ (3)
Finally, the metric is computed as the ratio of $\delta$ and $\mu$:
$\varepsilon=\frac{\delta}{\mu}.$ (4)
We refer to our metric as epsilon, as a tribute to the $\varepsilon$ symbol used in mathematics to denote error bounds. Algorithm 1 details the epsilon metric computation.
Algorithm 1 Algorithm for epsilon metric computation
  Select a batch of data from train set
  for arch in search space do
     Initialise empty output matrix
     for weight in [val1, val2do
        Initialise arch with constant shared weight
        Forward pass the batch through arch
        Get and flatten outputs
        Minmax normalise outputs (Eq. 1)
        Append outputs to the output matrix
     end for
     Compute difference between the rows of the output matrix (Eq. 2)
     Compute mean over the output matrix (Eq. 3)
     Compute epsilon metric (Equation 4)
  end for
2.1 Empirical evaluation
Here we evaluate the performance of epsilon and compare it to the results for zero-cost NAS metrics reported in Abdelfattah et al. (2021). We use the following evaluation scores (computed with NaN omitted):
  • Spearman $\rho$ (global): Spearman rank correlation $\rho$ evaluated on the entire dataset.
  • Spearman $\rho$ (top-$10\%$): Spearman rank correlation $\rho$ for the top-$10\%$ performing architectures.
  • Kendall $\tau$ (global): Kendall rank correlation coefficient $\tau$ evaluated on the entire dataset.
  • Kendall $\tau$ (top-$10\%$): Kendall rank correlation coefficient $\tau$ for the top-$10\%$ performing architectures.
[…]
  • Top-64/top-$10\%$: number of top-64 models ranked by zero-cost scoring metric within top-$5\%$ performing models.
2.1.1 NAS-Bench-201
The results for overall epsilon performance on NAS-Bench-201 are given in Table 1 along with other zero-cost NAS metrics. The Kendall $\tau$ score is not reported in Abdelfattah et al. (2021), but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation. We use the data provided by Abdelfattah et al. (2021) to evaluate their Kendal $\tau$. Note that our results differ from the original paper for some evaluation scores. In such cases, we indicate the original values between brackets. In particular, there is a discrepancy in computing the values in the last column, Top-64/top-$5\%$, while the rest of the results are consistent. Figure 6 in Appendix suggests that our calculations are correct.
Table 1: Zero-cost metrics performance for the NAS-Bench-201 search space with its three datasets: CIFAR-10, CIFAR-100 and ImageNet16-120. We give the original values from Abdelfattah et al. (2021) for reference between brackets. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
CIFAR-10
grad_sign 0.77
synflow 0.74 0.18 0.54 0.12 45.75 (46) 29 (44)
grad_norm 0.59 (0.58) -0.36 (-0.38) 0.43 -0.21 30.26 (30) 1 (0)
grasp 0.51 (0.48) -0.35 (-0.37) 0.36 -0.21 30.77 (30) 3 (0)
snip 0.60 (0.58) -0.36 (-0.38) 0.44 -0.21 30.65 (31) 1 (0)
fisher 0.36 -0.38 0.26 -0.24 4.99 ( 5) 0 (0)
epsilon 0.87 0.55 0.70 0.40 67.39 59
CIFAR-100
grad_sign 0.79
synflow 0.76 0.42 0.57 0.29 49.71 (50) 45 (54)
grad_norm 0.64 -0.09 0.47 -0.05 35.00 (35) 0 (4)
grasp 0.55 (0.54) -0.10 (-0.11) 0.39 -0.06 35.32 (34) 3 (4)
snip 0.64 (0.63) -0.08 (-0.09) 0.47 -0.05 35.25 (36) 0 (4)
fisher 0.39 -0.15 (-0.16) 0.28 -0.10 4.22 ( 4) 0 (0)
epsilon 0.90 0.59 0.72 0.43 81.24 62
ImageNet16-120
grad_sign 0.78
synflow 0.75 0.55 0.56 0.39 43.57 (44) 26 (56)
grad_norm 0.58 0.12 (0.13) 0.43 0.09 31.29 (31) 0 (13)
grasp 0.55 (0.56) 0.10 0.39 0.07 31.61 (32) 2 (14)
snip 0.58 0.13 0.43 0.09 31.16 (31) 0 (13)
fisher 0.33 0.02 0.25 0.01 4.61 ( 5) 0 (0)
epsilon 0.85 0.53 0.67 0.37 71.51 59
For NAS-Bench-201, we also report average performance when selecting one architecture from a pool of $N$ random architectures. The statistics are reported over $500$ runs. Table 2 compares epsilon to other trainless metrics. Note that te-nas starts with a super-network composed of all the edges and operators available within the space. In this case, $N$ is not applicable, and the performance can not be improved. In principle, other methods’ performance should improve with higher $N$ values.
Table 2: Comparison of the trainless metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. On the top, we list the best-performing methods that require training: REA (Real et al. (2019)), random search, REINFORCE (Williams (1992)), BOHB (Falkner et al. (2018)), DARTS- (Chu et al. (2020)). We report the average best-achieved test accuracy over $500$ runs, with $1{,}000$ architectures ($100$ for grad_sign) sampled from the search space at random. For tenas, the results are reported for $4$ random seeds. Random and optimal performances are given as a baseline.
Method Cost CIFAR-10 CIFAR-100 ImageNet16-120
(h) validation test validation test validation test
State-of-the-art
REA $3.3$ $91.19\pm 0.31$ $93.92\pm 0.3$ $71.81\pm 1.12$ $71.84\pm 0.99$ $45.15\pm 0.89$ $45.54\pm 1.03$
RS $3.3$ $90.93\pm 0.36$ $93.92\pm 0.31$ $70.93\pm 1.09$ $71.04\pm 1.07$ $44.45\pm 1.1$ $44.57\pm 1.25$
REINFORCE $3.3$ $91.09\pm 0.37$ $93.92\pm 0.32$ $71.61\pm 1.12$ $71.71\pm 1.09$ $45.05\pm 1.02$ $45.24\pm 1.18$
BOHB $3.3$ $90.82\pm 0.53$ $93.92\pm 0.33$ $70.74\pm 1.29$ $70.85\pm 1.28$ $44.26\pm 1.36$ $44.42\pm 1.49$
DARTS- $3.2$ $91.03\pm 0.44$ $93.80\pm 0.40$ $71.36\pm 1.51$ $71.53\pm 1.51$ $44.87\pm 1.46$ $45.12\pm 0.82$
Baselines (N=1000)
Optimal n/a $91.34\pm 0.18$ $94.20\pm 0.13$ $72.53\pm 0.53$ $72.84\pm 0.41$ $45.93\pm 0.51$ $46.59\pm 0.34$
Random n/a $84.11\pm 11.71$ $87.40\pm 11.94$ $61.57\pm 11.35$ $61.67\pm 11.35$ $33.97\pm 8.68$ $33.67\pm 8.98$
Trainless (N=1000)
naswot $3.3$ $89.69\pm 0.73$ $92.96\pm 0.81$ $69.86\pm 1.21$ $69.98\pm 1.22$ $43.95\pm 2.05$ $44.44\pm 2.10$
synflow n/a $89.91\pm 0.83$ $90.12\pm 0.78$ $70.35\pm 2.25$ $70.37\pm 2.08$ $41.73\pm 3.91$ $42.11\pm 4.02$
grad_norm n/a $88.13\pm 2.35$ $88.42\pm 2.28$ $66.35\pm 5.45$ $66.48\pm 5.32$ $33.88\pm 11.46$ $33.90\pm 11.74$
grasp n/a $87.85\pm 2.12$ $88.17\pm 2.04$ $65.36\pm 5.57$ $65.45\pm 5.48$ $32.23\pm 10.95$ $32.20\pm 11.23$
snip n/a $87.47\pm 2.19$ $87.81\pm 2.12$ $64.61\pm 5.52$ $64.74\pm 5.43$ $30.65\pm 11.32$ $30.55\pm 11.55$
fisher n/a $87.01\pm 2.31$ $87.36\pm 2.23$ $63.54\pm 5.69$ $63.67\pm 5.62$ $26.70\pm 10.83$ $29.56\pm 10.83$
epsilon $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
Trainless (N=100)
cv $0.05$ $84.89\pm 6.39$ $91.90\pm 2.27$ $63.99\pm 5.61$ $64.08\pm 5.63$ $38.68\pm 6.34$ $38.76\pm 6.62$
grad_sign n/a $89.84\pm 0.61$ $93.31\pm 0.47$ $70.22\pm 1.32$ $70.33\pm 1.28$ $42.07\pm 2.78$ $42.42\pm 2.81$
epsilon $0.003$ $\boldsymbol{90.44}\pm\boldsymbol{0.97}$ $\boldsymbol{93.39}\pm\boldsymbol{0.82}$ $\boldsymbol{70.85}\pm\boldsymbol{1.30}$ $\boldsymbol{71.00}\pm\boldsymbol{1.26}$ $\boldsymbol{44.03}\pm\boldsymbol{2.02}$ $\boldsymbol{44.20}\pm\boldsymbol{2.04}$
Trainless (all space)
tenas $0.5$ $93.9\pm 0.47$ $71.24\pm 0.56$ $42.38\pm 0.46$
Comparing the results for epsilon with other zero-cost NAS metrics, we can see that it outperforms them by a good margin. Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar figures for other methods can be found in Appendix, Figure 6). However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsilon development.
Refer to captionRefer to captionRefer to caption
Figure 1: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).
2.1.2 NAS-Bench-101
We use NAS-Bench-101 space to confirm that the success of epsilon metric in the previous section is not due to overfitting the NAS-Bench-201 search space and to see how it applies to a vaster search space. Table 3 together with Figure 2 confirm that it performs reasonably well on NAS-Bench-101, too.
Table 3: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. Values from Abdelfattah et al. (2021) are given for reference between brackets. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
grad_sign 0.45
synflow 0.37 0.14 0.25 0.10 22.67 (23) 4 (12)
grad_norm -0.20 -0.05 (0.05) -0.14 -0.03 1.98 (2) 0 (0)
grasp 0.45 -0.01 0.31 -0.01 25.60 (26) 0 (6)
snip -0.16 0.01 (-0.01) -0.11 0.00 3.43 (3) 0 (0)
fisher -0.26 -0.07 (0.07) -0.18 -0.05 2.65 (3) 0 (0)
epsilon 0.62 0.12 0.44 0.08 40.33 10
Refer to captionRefer to caption
Figure 2: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure shows $423{,}624$ and $14{,}322$ networks for NAS-Bench-101 and NAS-Bench-NLP, respectively (excluding architectures with NaN scores).
2.1.3 NAS-Bench-NLP
Both NAS-Bench-201 and NAS-Bench-101 are created to facilitate NAS in image recognition. They operate convolutional networks of very similar constitutions. To truly probe the generalisability of the epsilon metric, we test it on NAS-Bench-NLP. Both input data format and architecture type differ from the first two search spaces.
Unfortunately, Abdelfattah et al. (2021) provides no data for NAS-Bench-NLP, disabling us from using their results for calculations. Therefore, in Table 4, we give only values provided in the paper together with our epsilon metric (data for ficher is absent). We want to note that, unlike accuracy, perplexity used for language-related ML problems should be minimised. Therefore, the signs of correlations with scoring metrics should be reversed, which is not the case for numbers given in Abdelfattah et al. (2021).
Table 4: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
synflow 0.34 0.10 22
grad_norm -0.21 0.03 10
grasp 0.16 0.55 4
snip -0.19 -0.02 10
epsilon -0.34 -0.12 -0.23 -0.08 24.87 11
2.2 Integration with other NAS methods
While it is possible to utilise zero-cost metrics independently, they are often implemented within other NAS. Here we provide examples of random search and ageing evolution algorithms when used in tandem with epsilon.
Similarly to Abdelfattah et al. (2021), we compare random search performance with and without warm-up. First, we create a warm-up pool of $3{,}000$ architectures. Then, during the first $64$ steps of random search, the algorithm picks networks from the pool based on the highest trainless score and accordingly updates the best test performance.
For ageing evolution, the same principle applies. En plus, we report the results of implementation where every next parent is decided based on the highest epsilon score (”move”). In other words, the trainless scoring metric replaces validation accuracy in move mode. Finally, the child is created by parent mutation (within an edit distance of $1$) and added to the pool. Finally, the oldest network is removed from the pool.
For both described algorithms, we run the procedure until the number of trained architectures reaches $300$ and perform $100$ random rounds. Figure 3 shows that epsilon metric leads to considerable improvements in terms of time and precision. The best performance is achieved in combination with a warm-up. Figure 5 assembles warm-up performances for several trainless metrics.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: epsilon integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
3 Ablation studies
While our metric is relatively straightforward to implement, it does require several hyperparameters to be set. In this section, we provide the ablations studies’ results and analyse them.
3.1 Weights
To compute the epsilon metric, we initialise the networks with two distinct constant shared weights. The question is: how to fix their values, and how does this choice affect the whole method?
To answer this question, we ran a series of tests with multiple pairs of weights. In pairs, it makes sense to set the first weight always greater than the second one. This consideration stems from the fact that epsilon is based on the mean absolute difference between two initialisations. Therefore, it will output the same score for [$w_{1}$, $w_{2}$] and [$w_{2}$, $w_{1}$] combinations and will be exactly zero in the case where the two weights are equal. Tables 5, 6, 7 summarises the results of our tests. For the NAS-Bench-201 search space, the outcomes are very similar between the datasets, and the minor differences do not affect the conclusions drawn from this data.
Table 5: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
0.0001, 0.001 2120 0.47 0.41 0.33 0.28 38.21 23
0.0001, 0.01 2120 0.61 0.03 0.43 0.02 41.98 24
0.0001, 0.1 2120 0.61 0.03 0.43 0.02 41.98 24
0.0001, 1 2120 0.61 0.03 0.43 0.02 41.98 24
0.001, 0.01 4968 0.29 -0.05 0.20 -0.03 17.30 10
0.001, 0.1 4968 0.29 -0.05 0.19 -0.03 17.30 8
0.001, 1 4968 0.29 -0.05 0.19 -0.03 17.30 8
0.01, 0.1 5000 0.21 -0.12 0.14 -0.08 4.60 0
0.01, 1 5000 0.21 -0.12 0.14 -0.08 4.60 0
0.1, 1 5000 -0.47 0.04 -0.32 0.03 0.20 0
Table 6: Zero-cost metrics performance evaluated on NAS-Bench-201 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
1e-07, 1e-06 2957 0.33 -0.20 0.22 -0.13 4.76 1
1e-07, 1e-05 2957 0.39 -0.06 0.29 -0.04 9.15 2
1e-07, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-07, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-07, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-07, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-07, 1 2957 0.87 0.59 0.70 0.43 65.88 45
1e-06, 1e-05 2957 0.41 -0.06 0.30 -0.04 9.15 2
1e-06, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-06, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-06, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-06, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-06, 1 2957 0.87 0.59 0.70 0.43 65.88 45
1e-05, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-05, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-05, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-05, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-05, 1 2957 0.87 0.59 0.70 0.43 65.88 45
0.0001, 0.001 2957 0.62 0.06 0.43 0.05 19.26 3
0.0001, 0.01 2957 0.82 0.69 0.63 0.49 68.42 56
0.0001, 0.1 2957 0.84 0.64 0.65 0.46 68.14 56
0.0001, 1 2957 0.87 0.58 0.69 0.42 65.88 46
0.001, 0.01 4802 0.25 0.41 0.17 0.28 30.35 52
0.001, 0.1 4802 0.31 0.27 0.20 0.19 26.82 25
0.001, 1 4802 0.31 0.22 0.21 0.15 26.26 6
0.01, 0.1 4907 0.15 0.42 0.10 0.29 37.47 57
0.01, 1 4907 0.12 0.39 0.06 0.27 35.85 57
0.1, 1 4907 0.01 0.26 -0.03 0.16 0.84 0
Table 7: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
1e-07, 1e-06 2393 -0.24 0.32 -0.16 0.22 0.00 0
1e-07, 1e-05 2391 -0.22 0.52 -0.15 0.39 1.67 0
1e-07, 0.0001 2388 -0.35 0.49 -0.24 0.37 0.42 1
1e-07, 0.001 2332 -0.42 0.50 -0.29 0.37 2.99 1
1e-07, 0.01 1215 -0.30 0.13 -0.21 0.09 4.20 1
1e-07, 0.1 585 -0.49 0.55 -0.36 0.44 1.69 1
1e-07, 1 403 -0.12 0.08 -0.08 -0.00 12.20 7
1e-06, 1e-05 3286 -0.11 0.49 -0.07 0.36 0.91 0
1e-06, 0.0001 3283 -0.24 0.47 -0.16 0.34 0.62 0
1e-06, 0.001 3226 -0.33 0.44 -0.23 0.33 4.64 1
1e-06, 0.01 1639 -0.26 0.13 -0.18 0.09 3.40 2
1e-06, 0.1 841 -0.41 0.13 -0.28 0.12 10.71 1
1e-06, 1 594 -0.09 0.13 -0.06 0.08 11.67 7
1e-05, 0.0001 3948 -0.32 0.30 -0.21 0.22 0.51 0
1e-05, 0.001 3892 -0.35 0.20 -0.24 0.16 3.85 2
1e-05, 0.01 1900 -0.32 0.27 -0.22 0.18 2.86 1
1e-05, 0.1 913 -0.38 0.31 -0.28 0.25 13.04 1
1e-05, 1 649 -0.18 0.36 -0.12 0.25 10.77 7
0.0001, 0.001 3987 -0.35 0.18 -0.24 0.14 3.76 0
0.0001, 0.01 1943 -0.37 0.38 -0.26 0.25 3.59 1
0.0001, 0.1 925 -0.43 0.26 -0.30 0.21 11.83 1
0.0001, 1 663 -0.26 0.35 -0.18 0.24 10.45 7
0.001, 0.01 1936 -0.32 0.40 -0.21 0.26 2.59 3
0.001, 0.1 918 -0.38 0.27 -0.30 0.22 8.70 2
0.001, 1 663 -0.23 -0.20 -0.19 -0.17 7.46 5
0.01, 0.1 910 -0.36 0.04 -0.25 0.02 8.79 2
0.01, 1 654 -0.25 0.06 -0.16 0.04 9.09 5
0.1, 1 652 -0.27 0.20 -0.19 0.13 10.61 7
The results show that there is indeed some variation between different initialisations. Two effects impact performance. Firstly, the more significant the difference between the two weights, the higher the correlation between epsilon and trained performance. It makes sense since if the weights are too close to each other, the difference between the outputs becomes minimal, and the distinction between architectures in terms of epsilon is no longer possible.
Secondly, we can see that too low and too high weights result in fewer architectures with non-NaN scores. Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsilon metric. Naturally, the deeper the network is, the more signal attenuation is, and the higher the probability of NaN score. We remind that prior to epsilon computation, the outputs are normalised to the same range $[0,1]$, which means the value of the metric per se does not become smaller for deeper networks. This extinction effect is due to the limited sensitivity of float type. Theoretically, with infinite float sensitivity, there would be no extinction and hence no effects on the choice of the weights.
Table 8: Optimal weights for epsilon evaluation for three search spaces and their datasets.
Search space Dataset Optimal weights
NAS-Bench-101 CIFAR-10 $[10^{-4},10]$
NAS-Bench-201 CIFAR-10 $[10^{-7},1]$
CIFAR-100 $[10^{-7},1]$
ImageNet16-120 $[10^{-7},1]$
NAS-Bench-NLP PTB $[10^{-5},10^{-3}]$
As a rule of thumb, for every new ML problem, we suggest running epsilon evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs. While we acknowledge that this procedure is subjective, importantly, it does not require any training.
Weight combinations that we selected for each search space are given in Table 8.
3.2 Initialisation algorithm
The choice to initialise the weights with constant values provides a ground for discussion. On the one hand, constant shared weights guarantee the same initialisation regardless of the structure; the noise from the randomness of other initialisations is not obstructing the effect of the neural geometry. It would be possible to reduce the amount of noise by increasing the number of independent initialisations with distinct random seeds similar to Gracheva (2021), but this increases the algorithm run-time proportionally.
Table 9 gives the results of the ablation studies when we compute epsilon based on two initialisations.
Table 9: Initialisations ablation studies. The metric is computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
Initialisation Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Uniform 15284 0.14 -0.17 0.09 -0.11 3.07 7
Normal 15284 0.44 -0.29 0.30 -0.20 20.07 4
Kaiming uniform 15284 0.32 0.11 0.21 0.07 27.40 26
Kaiming normal 15284 0.49 -0.29 0.33 -0.20 17.86 7
Orthogonal 15284 0.41 -0.20 0.28 -0.13 15.11 1
3.3 Batch size
To test the method’s sensitivity against the batch size, we ran the scoring routine with $8$ different batch sizes $[8,16,32,64,128,512,1024]$, each with $10$ random batches. The tests are done on the NAS-Bench-201 search space with its three datasets. For each batch, we report medians with $25$ and $75$ percentiles over $10$ runs. We evaluate epsilon on $5{,}000$ architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Batch size ablation study for NAS-Bench-201 search space, CIFAR-10 (left), CIFAR-100 (centre) and ImageNet16-120 (right) datasets.
Figure 4 show that, as expected, the larger the batch size, the better and more stable the performance of epsilon metric. There is almost no improvement for batches over $256$. We set the batch size to this value in the experiments throughout the paper.
3.4 Embedding initialisation
In order to verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
  • uniform positive 0.1: random uniform from ranges [0, 0.1]
  • uniform positive 1: random uniform from ranges [0, 1]
  • uniform centred 0.1: random uniform from ranges [-0.1, 0.1]
  • uniform centred 1: random uniform from ranges [-1, 1]
  • random 0.1: random normal centred at 00 with a standard deviation of 0.1
  • random 1: random normal centred at 00 with a standard deviation of 1
Table 10 summarises our results of the embedding ablations. Ablation is done on the NAS-Bench-NLP search space, PTB dataset (the only search space implementing embedding).
Table 10: Embedding ablation studies. Metric computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
Embedding Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Uniform positive 0.1 782 -0.38 0.17 -0.26 0.15 3.80 2
Uniform positive 1 776 -0.37 0.21 -0.26 0.18 2.56 2
Uniform centered 0.1 783 -0.40 0.18 -0.28 0.15 2.53 2
Uniform centered 1 782 -0.46 0.18 -0.33 0.16 1.27 1
Random 0.1 783 -0.42 0.18 -0.29 0.15 2.53 2
Random 1 782 -0.47 0.19 -0.33 0.16 1.27 1
Care should be taken when initialising the networks containing the embedding layer: if embedding is initialised with all constants, there is no difference between the embedded input. The performance of our metric, in this case, will be analogous to that with a batch size of one.
Our ablation studies show that as soon as the embedding is initialised with non-constant weights, it does not significantly influence the outcomes.
3.5 Synthetic data
In this section, we test the importance of the input data by feeding our metric several synthetic input data on CIFAR-10. We feed networks with $5$ types of data:
  • actual data: a batch of CIFAR-10 data
  • grey scale images: images within the batch are solid colour ranging from black to white
  • random normal: images are filled with random values following normal distribution with $[\mu,\sigma]=[0,1]$
  • random uniform: images are filled with random values following uniform distribution with $[\mu,\sigma]=[-1,1]$
  • random uniform (+): images are filled with random values following uniform distribution with $[min,\max]=[0,1]$
All the tests are performed with batch size of $256$, weights $[10^{-7},1]$ and $5{,}000$ architectures. Table 11 shows that even though the performance with synthetic data drops compared to actual data, it is still reasonably good.
Curiously, greyscale images, filled with constant values, show the closest results to the actual data. Note that epsilon metric with greyscale data outperforms synflow. It is an essential achievement because synflow does not use input data and is, therefore, data independent. Our results show that epsilon has the potential to be used with no data whatsoever.
Table 11: Synthetic data tests with CIFAR-10 dataset, NAS-Bench-201 search space. Tests are based on the evaluation of $5{,}000$ architectures.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Real data 0.87 0.59 0.70 0.43 65.88 $45$
Grey scale 0.87 0.44 0.68 0.31 59.46 $34$
Random normal 0.54 0.15 0.38 0.14 14.97 $3$
Random uniform 0.56 0.17 0.40 0.16 14.19 $3$
Random uniform (+) 0.61 0.17 0.43 0.16 16.22 $4$
4 Visualisation of other zero-cost NAS metrics
4.1 When integrated with other NAS methods
In Section 2.2, we show how epsilon metric improves the performance of ageing evolution and random search when used for warming up. Figure 5 compares epsilon integration to other metrics from Abdelfattah et al. (2021). For both EA ans RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds. The warm-up pool contains $3000$ randomly selected architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Various zero-cost NAS metrics’ performance when used as a warm-up for ageing evolution (top) and random search (bottom) for three datasets from the NAS-Bench-201 search space. Shadows span between lower and upper quartiles based on $100$ rounds. The warm-up population is $3000$ architectures.
4.2 Correlation with accuracy
In Abdelfattah et al. (2021) the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding. Here we provide visualisations for two search spaces built on the data provided by the authors.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from Abdelfattah et al. (2021). Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. Figure represents search space of $15{,}625$ networks (excluding architectures with NaN scores).
5 Discussion
The epsilon metric performs surprisingly well on various benchmark datasets. Similarly to other NAS algorithms, it shows the most robust performance with the NAS-Bench-201 benchmark. This fact indicates that the set of architectures comprising NAS-Bench-201 forms a relatively easy-to-navigate landscape regardless of the metric. Still, epsilon shows the best performance among trainless metrics. Table 1 shows that synflow shows comparable performance in terms of correlations on the ImageNet-16-120 dataset. However, statistically epsilon selects, on average, a better-performing architecture, which can also be confirmed visually by comparing Figures 1 and 6.
Notably, we would like to highlight the difference with our previous work (cv metric, Gracheva (2021)). The two methods differ in how the weights are de-emphasised and what outputs we use for statistics computation. Two constant shared weight initialisations better bring out the effect of the bare geometry compared to a hundred random initialisations. Moreover, epsilon metric takes significantly less time to compute, being fifty-fold faster.
It is important to note that we used NAS-Bench-201 for the metric development; therefore, the results of this part of the study might be biased.
To confirm the generalisability of the metric, we applied it to the NAS-Bench-101 benchmark. Again, Table 3 shows strong performance. The peculiar vertical lines are due to clustering in terms of test accuracy. Since every architecture in NAS-Bench-101 is trained with $3$ random seeds, the four clusters correspond to situations when the training diverges for some random initialisations. More random seeds would reveal a smoother picture and better correlations. Again, we observe comparable performance between synflow and epsilon scoring metrics. We could only confirm some of the results from Abdelfattah et al. (2021) and suppose that some statistics might have been miscalculated.
By design, we expect our metric to be independent of the specific neural architecture structure. To confirm this claim, this is essential to show its applicability to architectures other than CNNs. Our work tests epsilon applicability to RNN models with NAS-Bench-NLP.
While statistics in Table 4 are not exceptional, Figure 2 shows a visible trend towards better architectures with increasing metric value. Significant noise levels might stem from the characteristics of the benchmark itself: factors like a relatively small sample of networks given a vast space, choice of training hyperparameters, dropout rates, and others may distort NAS performance. Nonetheless, the trend is strong enough to conclude that epsilon metric can apply to recurrent type architectures.
While epsilon metric shows solid empirical performance, the underlying reasons for this are unclear.
There are several hints towards its understanding. First, mathematically, epsilon represents the difference in the output distribution shapes between initialisations. The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry. By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
Second, according to the weight ablation studies (Section 3), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive outputs explosion or vanishing. Therefore, epsilon measures the amplitude of the outputs’ distribution shape change due to geometry.
Finally, we see that grey-scale solid images work reasonably well as inputs during the synthetic data studies. The distribution over the input samples is uniform, which makes it easier to track the changes in the distribution shape as the signal propagates through the network.
6 Conclusion
This work presents a novel zero-cost NAS scoring metric epsilon. It allows evaluating a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. The value of epsilon reflects how neural network outputs’ distribution changes between low and high constant shared weight initialisations. It shows that the higher this difference, the better the network will perform upon training.
It takes a fraction of a second to evaluate a single architecture and does not require gradients computation (low demand on memory) or labels, and can run on a CPU. We provide experimental evidence that the introduced epsilon metric generalises to various architecture types, such as convolutional or recurrent neural networks (care should be taken with embedding layer initialisation).
This work evaluates epsilon on three staple NAS search spaces: NAS-Bench-201, NAS-Bench-101 and NAS-Bench-NLP. It sets the state-of-the-art among zero-cost NAS metrics and can be easily implemented within other methods, such as random search or evolutionary algorithms.
The only significant disadvantage of the method is that it requires the choice of constant weight values used for initialisation. Our tests show that it must be set up individually for each search space. We plan to automate the weight selection process in our future work.
Overall, while a coherent theoretical foundation of epsilon is missing, we consider the empirical results compelling.

### v3

@@
-Neural Architecture Search via Two Constant Shared Weights Initialisations
+Neural Architecture Search: Two Constant Shared Weights Initialisations
 Abstract
-In recent years, zero-cost metrics are gaining ground in neural architecture search (NAS).
-These metrics allow finding the optimal neural network for a given task faster and with a lesser computational load than conventional NAS methods.
-Equally important is that they also shed some light on the internal workings of neural architectures.
-This paper presents a zero-cost metric that highly correlates with the train set accuracy across the NAS-Bench-101, NAS-Bench-201 and NAS-Bench-NLP benchmark datasets.
-We evaluate a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations.
-For this, we only use an unlabelled mini-batch of data.
-We observe that the dispersion of the outputs between two initialisations positively correlates with trained accuracy.
-The correlation further improves when we normalise dispersion by average output magnitude.
-The resulting metric, epsilon, does not require gradients computation and unbinds the NAS procedure from training hyperparameters, loss metrics and human-labelled data.
-Our method is easy to integrate within existing NAS algorithms and takes a fraction of a second to evaluate a single network.
+In the last decade, zero-cost metrics have gained prominence in neural architecture search (NAS) due to their ability to evaluate architectures without training.
+These metrics are significantly faster and less computationally expensive than traditional NAS methods and provide insights into neural architectures’ internal workings.
+This paper introduces epsinas, a novel zero-cost NAS metric that assesses architecture potential using two constant shared weight initialisations and the statistics of their outputs.
+We show that the dispersion of raw outputs, normalised by their average magnitude, strongly correlates with trained accuracy.
+This effect holds across image classification and language tasks on NAS-Bench-101, NAS-Bench-201, and NAS-Bench-NLP.
+Our method requires no data labels, operates on a single minibatch, and eliminates the need for gradient computation, making it independent of training hyperparameters, loss metrics, and human annotations.
+It evaluates a network in a fraction of a GPU second and integrates seamlessly into existing NAS frameworks.
 The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
-Keywords: machine learning, neural architecture search, zero-cost NAS
+keywords:
+machine learning, neural architecture search, zero-cost NAS, efficient NAS, NAS benchmarks
 1 Introduction
-The field of neural architecture search (NAS) emerged about a decade ago as an effort to automatise the process of neural geometry optimisation.
-At the early stages of NAS development, every candidate architecture used to be evaluated through the training process (reinforcement learning (Williams (1992)), evolutionary algorithms (Real et al.
-(2019)), Bayesian optimisation (Falkner et al.
-(2018); White et al.
-(2021))).
-One-shot algorithms adopting weight sharing dispense of multiple architectures training, reducing the search time drastically (efficient reinforcement learning (Pham et al.
-(2018)), random search with parameters sharing (Li and Talwalkar (2020)), differentiable methods).
-Nevertheless, they require the training of a massive hypernet, which necessitates elaborate hyperparameter tuning.
-While these methods prove efficient, they do not systematically achieve satisfactory results (Dong and Yang (2019)).
-One of the best of them, DARTS- (Chu et al.
-(2020)), shows significant uncertainty compared to evolutionary or reinforcement algorithms.
-There are available methods that estimate network performance without training the dataset of interest but relying on an auxiliary predictive machine learning (ML) model built on a dataset of trained architectures (Istrate et al.
-(2019); Deng et al.
-(2017)).
-These methods aim to accelerate the NAS process for image recognition but still rely on training and cannot apply to other ML problems.
+The field of neural architecture search (NAS) emerged roughly a decade ago as a solution to automate the process of neural geometry optimisation.
+In its early stages, NAS predominantly relied on evaluating candidate architectures through computationally intensive training processes, using techniques such as reinforcement learning [49], evolutionary algorithms [38, 31], and Bayesian optimisation [13, 47].
+One-shot algorithms speed up the search by sharing weights instead of training each architecture separately.
+These methods include efficient reinforcement learning [37], random search with parameter sharing [27], and differentiable approaches [30, 7, 50].
+However, they require training a large hypernetwork, necessitating intricate hyperparameter tuning.
+While these techniques are efficient, they often fail to deliver consistent results [11].
+For example, even advanced approaches like DARTS- [7] exhibit significant variability compared to evolutionary or reinforcement learning methods.
+Some methods attempt to estimate network performance without training on the dataset of interest by leveraging an auxiliary predictive machine learning (ML) model built on a dataset of pre-evaluated architectures [19, 9].
+Although these methods accelerate the NAS process for image recognition, they still rely on training and are not applicable to other domains of ML.
 […]
-1.1 Zero-cost NAS
-To alleviate the process of architecture search, many researchers focus on developing methods that allow finding optimal architectures without model training – so-called zero-cost NAS methods.
-These methods evaluate networks via a trainless proxy.
-They typically require the equivalent of one or a few training epochs, which is two to three orders of magnitude faster than previous NAS methods.
-Weight agnostic neural networks.
-One of the pioneering works in zero-shot NAS is presented by Gaier and Ha (2019).
-The authors demonstrate a constructor building up neural architectures based on the mean accuracy over several initialisations with constant shared weights and the number of parameters contained within the model.
-The resulting model achieves over $90\%$ accuracy on MNIST data (LeCun et al.
-(2010)) when the weights are fixed to the best-performing constants.
-While these results are very intriguing, the authors admit that such architectures do not perform particularly well upon training.
-Moreover, back in $2019$, the benchmark databases of trained architectures, which are now routinely used to compare NAS metrics with each other, were yet to be released, which disables the comparison of this zero-shot method against the most recent ones.
+2 Related work
+To simplify architecture search, zero-cost NAS methods identify optimal architectures without requiring full training.
+These methods leverage zero-cost proxies [16, 34, 14, 18, 25], typically demanding computational resources equivalent to just one or a few training epochs.
+Consequently, zero-cost NAS is two to three orders of magnitude faster than traditional NAS approaches.
+Below, we summarize the most prominent zero-shot NAS methods.
+For a comprehensive field review, see [26].
+[…]
 ReLU activation pattern.
-In $2020$, Mellor et al.
-(2020) present the naswot metric, which exploits the rectified linear unit (ReLU, Agarap (2018)) activation function’s property to yield distinct activation patterns for different architectures.
-Concretely, every image yields a binary activation vector upon passing through a network, forming a binary matrix for a mini-batch.
+In 2020, [34] introduced the naswot metric, which exploits the property of the rectified linear unit (ReLU) activation function [2] to produce distinct activation patterns for different architectures.
+Specifically, every image passing through a network generates a binary activation vector, which forms a binary matrix for a mini-batch.
 The logarithm of the determinant of this matrix serves as a scoring metric.
-Authors show that larger naswot values are associated with better training performances, which leads to the conclusion that high-performing networks should be able to distinguish the inputs before training.
-Unfortunately, the method can only be implemented on networks with ReLU activation functions, which limits its applicability to convolutional architectures.
-Another work employing the aforesaid ReLU property is the one from Chen et al.
-(2021).
-They combine the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to build the tenas metric.
-Instead of evaluating each network in the search space individually, they create a super-network built with all the available edges and operators and then prune it.
+The authors demonstrated that higher naswot values are associated with better training performance, concluding that high-performing networks should be able to distinguish inputs even before training.
+However, this method is limited to networks using ReLU activation functions, which restricts its applicability to convolutional architectures.
+A conceptually similar work by [4] combines the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to develop the tenas metric.
+Instead of individually evaluating each network in the search space, they construct a super-network that includes all the available edges and operators, which is then pruned.
 Coefficient of variance.
-Our previous work on fully trainless NAS (Gracheva (2021)) evaluates the stability of untrained scores over random weights initialisations.
-The networks are initialised with multiple random seeds, and architectures are selected based on the coefficient of variance of the accuracy at initialisation, cv.
-While cv performance is associated with a high error rate, the author concludes that a good architecture should be stable against random weight fluctuations.
-While this method can, in theory, apply to any neural architecture type, it requires multiple initialisations and is relatively heavy compared to naswot and later methods.
-Furthermore, accuracy-based scoring metrics can only apply to classification problems, and it is unclear how to extend cv implementation to the regression tasks.
+Our previous work on fully trainless NAS [18] evaluated the stability of untrained scores across random weight initialisations.
+Networks were initialised with multiple random seeds, and architectures were selected based on the coefficient of variance of the accuracy at initialisation, cv.
+While cv performance is associated with a high error rate, the study concluded that a good architecture should demonstrate stability against random weight fluctuations.
+Although this method theoretically applies to any type of neural architecture, it requires multiple initialisations and is relatively computationally expensive compared to naswot and later methods.
+Additionally, accuracy-based scoring metrics are limited to classification problems, and it remains unclear how to adapt the cv approach for regression tasks.
+Inverse variance of gradients.
+A recent study by [25] introduced the zico metric, a zero-cost proxy for neural architecture performance that, like cv [18], is derived from network statistics.
+zico is computed as the ratio of the average gradient values to their standard deviation—in essence, a reversed cv, where accuracy is substituted with gradient values.
+Notably, zico exhibits a strong correlation with trained accuracy, surpassing nparams—the total number of parameters in a network—as a performance estimator.
+While nparams may seem like a simplistic proxy, it remained the most effective predictor of network performance until 2023.
+Gradient signal-to-noise ratio.
+In 2023, [44] introduced the gradient signal-to-noise ratio metric ($\xi$-gsnr), which, as the name suggests, closely resembles zico.
+Their approach builds on zico by squaring the metric and incorporating a small constant $\xi$ in the denominator.
+This seemingly minor modification significantly impacts the metric’s performance, establishing a new state-of-the-art zero-cost NAS.
+Additionally, the authors provide essential theoretical justification for their approach.
+A comparison of the cv, zico, and $\xi$-gsnr methodologies is presented in Table 1.
 Gradient sign.
-The grad_sign metric is built to approximate the sample-wise optimisation landscape (Zhang and Jia (2021)).
-The authors argue that the closer local minima for various samples sit to each other, the higher the probability that the corresponding gradients will have the same sign.
-The number of samples yielding the same gradient sign approximates this probability.
-It allows to evaluate the smoothness of the optimisation landscape and architecture trainability.
-The method requires labels and gradient computation.
+The grad_sign metric is designed to approximate the sample-wise optimisation landscape [52].
+The authors argue that when local minima for various samples are closer, the probability increases that the corresponding gradients will have the same sign.
+The number of samples yielding the same gradient sign approximates this probability, allowing for the smoothness evaluation of the optimisation landscape and the architecture’s trainability.
+However, this method requires labels and gradient computation.
 Pruning-at-initialisation proxies.
-Several powerful zero-cost proxies have emerged as an adaptation of pruning-at-initialisation methods to NAS in the work by Abdelfattah et al.
-(2021): grad_norm (Wang et al.
-(2020)), snip (Lee et al.
-(2018)), synflow (Tanaka et al.
-(2020)).
-These metrics are originally developed to evaluate the network’s parameters’ salience and prune away potentially meaningless synapses.
-They require a single forward-backwards pass to compute the loss.
-Then, the importance of parameters is computed as a multiplication of the weight value and gradient value.
-Then, they integrate the salience over all the parameters in the network to evaluate its potential upon training.
-What is particularly interesting about the synflow metric is that it evaluates the architectures without looking at the data by computing the loss as the product of all the weights’ values (randomly initialised).
-synflow metric shows the most consistent performance among various search spaces and sets the state-of-the-art for the zero-cost NAS.
-Both naswot and synflow do not depend on labels, which arguably reduces the effect of human error during data labelling.
+Several promising zero-cost proxies have emerged as adaptations of pruning-at-initialisation methods for NAS, as demonstrated in the work by [1]: grad_norm [46], snip [24], and synflow [45].
+These metrics were originally developed to assess the salience of a network’s parameters and prune potentially irrelevant synapses.
+They require a single forward-backwards pass to compute the loss, after which the importance of parameters is calculated as the product of the weight and gradient values.
+The salience is then integrated over all parameters in the network to estimate its potential performance once trained.
+What makes the synflow metric particularly notable is its ability to evaluate architectures without referencing the data, as it computes the loss based solely on the product of all randomly initialised weights’ values.
+Among these proxies, the synflow metric demonstrates the most consistent performance across various search spaces and sets the state-of-the-art for zero-cost NAS.
+Neither naswot nor synflow depend on labels, which reduces the effect of human error during data labelling.
 Moreover, naswot does not require gradient computation, which renders this method less memory-intensive.
-The above results imply that neural networks might have some intrinsic property which defines their prediction potential before training.
-Such property should not depend on the values of trainable parameters (weights) but only on the network’s topology.
-In the present work, we combine the takeaways from the existing trainless NAS implementations to present a new metric which significantly outperforms existing zero-cost NAS methods.
-1.2 NAS benchmarks
-To guarantee our metric’s reproducibility and compare its performance against other NAS algorithms, we evaluate it on the three widely used NAS benchmark datasets.
-NAS-Bench-101.
-The first and one of the largest NAS benchmark sets of trained architectures.
-It consists of $423{,}624$ cell-based convolutional neural networks (Ying et al.
-(2019)).
-The architectures consist of three stacks of cells, each followed by max-pooling layers.
-Cells may have up to $7$ vertices and $9$ edges, with $3$ possible operations.
-This benchmark is trained multiple times on a single dataset, CIFAR-10 (Krizhevsky et al.
-(2009)), with a fixed set of hyperparameters for $108$ epochs.
-NAS-Bench-201.
-It is a set of architectures with a fixed skeleton consisting of a convolution layer and three stacks of cells connected by a residual block (Dong and Yang (2020)).
-Each cell is a densely connected directed acyclic graph with $4$ nodes, $5$ possible operations and no limits on the number of edges, providing a total of $15{,}625$ possible architectures.
-Architectures are trained on three major datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al.
-(2009)) and a downsampled version of ImageNet (Chrabaszcz et al.
-(2017)).
-Training hyperparameters are fixed, and the training spans $200$ epochs.
-NAS-Bench-NLP.
-As the name suggests, this benchmark consists of architectures suitable for neural language processing (Klyuchnikov et al.
-(2022)).
-Concretely, it consists of randomly-generated recurrent neural networks.
-Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations.
-Here, we only consider models trained and evaluated on Penn Treebank (PTB, Marcinkiewicz (1994)) dataset: $14{,}322$ random networks with a single seed and $4{,}114$ with tree seeds.
-The training spans $50$ epochs and is conducted with fixed hyperparameters.
-2 Epsilon metric
-Two existing NAS methods inspire the metric that we share in the present work: cv (Gracheva (2021)) and weight agnostic neural networks (Gaier and Ha (2019)).
-Both metrics aim to exclude individual weight values from consideration when evaluating networks: the former cancels out the individual weights via multiple random initialisations, while the latter sets them to the same value across the network.
-It is very intriguing to see that an network can be characterised purely by its topology.
-As mentioned above, the cv metric has two principal disadvantages.
-While it shows a fairly consistent trend with trained accuracy, it suffers from high uncertainty.
-It can be, to some degree, explained by the fact that random weight initialisations bring in some noise.
-Our idea is that replacing random initialisations with single shared weight initialisations, similarly to Gaier and Ha (2019), should improve the method’s performance.
-The second weak point is that cv is developed for classification problems and relies on accuracy.
-Therefore, it needs to be clarified how to apply this metric to regression problems.
-The coefficient of variation is a ratio of standard deviation to mean, and Gracheva (2021) shows that cv correlates negatively with train accuracy.
-It implies that the mean untrained accuracy should be maximised.
-On the other hand, for regression tasks, performance is typically computed as some error, which is sought to be minimised.
+A few other efficient zero-shot methods cannot be directly compared to our results (such as Zen-NAS by [28]).
+For more details on the current status of NAS, please refer to the most recent systematic review on NAS [40].
+The results of the above studies suggest that neural networks possess an intrinsic property that determines their predictive potential even before training.
+This property should be independent of the specific values of trainable parameters (weights) and instead depend solely on the network’s topology.
+In this work, we build upon insights from existing zero-cost NAS implementations to propose a new metric that shows strong performance and outperforms current zero-cost NAS methods.
+3 Epsinas metric
+Two existing NAS methods inspire the metric that we share in the present work: cv [18] and weight agnostic neural networks [16].
+Both metrics aim to reduce the role of individual weights in the network to access its topology.
+The cv metric levels out the individual weights via multiple random initialisations, while [16] sets the weights to the same value across the network.
+As mentioned, the cv metric has two principal disadvantages.
+While it shows a fairly consistent trend with trained accuracy, it suffers from high statistical deviations.
+To some degree, it must be due to the noise from random weight initialisations.
+With epsinas, we replaced random initialisations with single shared weight initialisations to improve the method’s performance.
+The second weak point of cv is that it has been developed for classification problems and relies on accuracy, which should be maximised.
+On the other hand, for regression tasks, performance is typically computed as an error, which is sought to be minimised.
+The cv is a ratio of standard deviation and the mean of untrained accuracies, and the final metric correlates negatively with train accuracy.
 It is not apparent whether the division by mean untrained error would result in the same trend for cv metric.
-To address this issue, we decided to consider raw outputs.
+To address this issue, we decided to step out from evaluation metrics and consider raw outputs instead.
 This modification renders the method applicable to any neural architecture.
-However, it comes with a difference: accuracy returns a single value per batch of data, while raw outputs are $[N_{\textrm{BS}}\times L]$ matrices, where $N_{\rm{BS}}$ is the batch size and $L$ is the length of a single output.
-[^1] This length depends on the task and architecture: for regression tasks $L=1$, for classification, $L$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string.
-In our work, we flatten these matrices to obtain a single vector $\mathbf{v}$ of length $L_{v}=N_{\textrm{BS}}\times L$ per initialisation.
-We then stuck both initialisations into a single output matrix $\mathbf{V}$.
+For this, we flatten the raw output matrices to obtain a single vector $\mathbf{v}$ of length $L_{\mathbf{v}}=N_{\textrm{BS}}\times L_{\mathbf{O}}$ per initialisation, where $N_{\rm{BS}}$ is the batch size and $L_{\mathbf{O}}$ is the length of a single output  [^1] This length depends on the task and architecture: for regression tasks $L_{\mathbf{O}}=1$, for classification, $L_{\mathbf{O}}$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string..
 Before proceeding to statistics computation over initialisations, we also must normalise the output vectors: in the case of constant shared weights, outputs scale with weight values.
-In order to compare initialisations on par with each other, we use min-max normalisation:
+To compare initialisations on par with each other, we use min-max normalisation:
 $\mathbf{v}_{i}^{\prime}=\frac{\mathbf{v}_{i}-\min(\mathbf{v}_{i})}{\max(\mathbf{v}_{i})-\min(\mathbf{v}_{i})},$ (1)
@@
 Accordingly, instead of standard deviation, we use mean absolute error between the normalised outputs of two initialisations:
-$\delta=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\lvert\mathbf{V}_{1,j}^{\prime}-\mathbf{V}_{2,j}^{\prime}\rvert.$ (2)
+$\mathrm{MAE}=\frac{1}{L_{\mathbf{v}}}\sum\limits_{j=0}^{L_{\mathbf{v}}}\lvert\mathbf{v}_{0,j}^{\prime}-\mathbf{v}_{1,j}^{\prime}\rvert.$ (2)
 The mean is computed over the outputs of both initialisations as follows:
-$\mu=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\frac{\mathbf{V}_{1,j}^{\prime}+\mathbf{V}_{2,j}^{\prime}}{2}=\frac{1}{2L_{v}}\sum\limits_{i=0}^{2}\sum\limits_{j=0}^{L_{v}}\mathbf{V}_{i,j}^{\prime}$ (3)
-Finally, the metric is computed as the ratio of $\delta$ and $\mu$:
-$\varepsilon=\frac{\delta}{\mu}.$ (4)
-We refer to our metric as epsilon, as a tribute to the $\varepsilon$ symbol used in mathematics to denote error bounds. Algorithm 1 details the epsilon metric computation.
-Algorithm 1 Algorithm for epsilon metric computation
+$\mu=\frac{1}{L_{\mathbf{v}}}\sum\limits_{j=0}^{L_{\mathbf{v}}}\frac{\mathbf{v}_{1,j}^{\prime}+\mathbf{v}_{2,j}^{\prime}}{2}$ (3)
+Finally, the metric is computed as the ratio of $\mathrm{MAE}$ and $\mu$:
+$\varepsilon=\frac{\mathrm{MAE}}{\mu}.$ (4)
+We use real data from the respective datasets to compute the statistics, as suggested in [14]. In Section 6.5, we demonstrate that using real data enhances search efficiency.
+We refer to our metric as epsinas, paying tribute to the $\varepsilon$ symbol commonly used in mathematics to denote error bounds.
+The epsinas metric was developed independently but in parallel with zico and follows a similar approach.
+While zico uses the mean and standard deviation of gradients over two mini-batches, our metric combines the mean and mean absolute error of raw network outputs.
+This parallel development highlights a shared idea — leveraging statistical properties of untrained networks to guide architecture selection.
+Table 1 compares epsinas to the three the most resembling zero-cost proxies: cv, zico, and $\xi$-gsnr.
+Although the difference between the two methods may seem minor and zico and $\xi$-gsnr show a strong correlation with trained accuracy, epsinas has an advantage: it does not require backpropagation for gradients computation, making it faster to calculate.
+Additionally, as we demonstrate in the next section, epsinas exhibits an even stronger correlation with trained accuracy.
+Table 1: Comparison of epsinas to the three the most resembling zero-cost proxies.
+Data $\mathbf{D}$ denotes the data used for metrics computation.
+$N_{init}$ refers to the number of initialisations (for zico and $\xi$-gsnr we report the number of batches).
+Metric Data, $\mathbf{D}$ Formula Initialisation $N_{init}$ Gradients True labels
+cv Accuracy $\frac{\sigma(\mathbf{D})}{\mu(\mathbf{D})}$ Kaiming normal 100 No Yes
+zico Gradients $\frac{\mu(\mathbf{D}}{\sigma(\mathbf{D})}$ Kaiming normal 2 Yes Yes
+$\xi$-gsnr Gradients $\frac{\mu(\mathbf{D})^{2}}{\sigma(\mathbf{D})^{2}+\xi}$ Kaiming normal 8 Yes Yes
+epsinas Normalised outputs $\frac{\mathrm{MAE}(\mathbf{D})}{\mu(\mathbf{D})}$ Constant shared weights 2 No No
+Algorithm 1 details the epsinas metric computation.
+Algorithm 1 Algorithm for epsinas metric computation
   Select a batch of data from train set
@@
      Compute mean over the output matrix (Eq. 3)
-     Compute epsilon metric (Equation 4)
+     Compute epsinas metric (Equation 4)
   end for
-2.1 Empirical evaluation
-Here we evaluate the performance of epsilon and compare it to the results for zero-cost NAS metrics reported in Abdelfattah et al.
-(2021).
-We use the following evaluation scores (computed with NaN omitted):
+Proper initialization is crucial for implementing the metric effectively.
+In our current implementation, we preserve the initialization of biases, batch normalization, and embeddings.
+Specifically, bias weights are set to zero, batch normalization parameters are initialized to ones, and embeddings follow the original paper’s approach using random integers [20].
+4 Experimental design
+4.1 NAS benchmarks
+Following standard NAS metric evaluation practices, we assess epsinas on three well-established NAS benchmark datasets containing fully trained neural architectures.
+To ensure the broad applicability of epsinas across different architectures and tasks, we evaluate it on both image classification and language processing tasks.
+NAS-Bench-101.
+The first and one of the most extensive NAS benchmarks, NAS-Bench-101, consists of $423{,}624$ fully trained convolutional neural networks [51].
+Each architecture comprises three stacked cells, followed by max-pooling layers, with up to $7$ vertices, $9$ edges, and $3$ possible operations per cell.
+All models are trained on CIFAR-10 [21] for $108$ epochs using fixed hyperparameters.
+NAS-Bench-201.
+This benchmark provides $15{,}625$ architectures with a fixed backbone: a convolutional layer followed by three stacked cells connected by a residual block [12].
+Each cell is a densely connected directed acyclic graph with $4$ nodes and $5$ possible operations.
+Architectures are trained on CIFAR-10, CIFAR-100 [21], and a downsampled version of ImageNet [5], all with fixed hyperparameters over $200$ epochs.
+NAS-Bench-NLP.
+Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations.
+Here, we only consider models trained and evaluated on Penn Treebank (PTB) data [33]: $14{,}322$ random networks with a single seed.
+The training spans $50$ epochs and is conducted with fixed hyperparameters.
+4.2 Evaluation metrics
+We evaluate the performance of epsinas and compare it to the results for zero-cost NAS metrics reported in [1]. We use the following evaluation scores (computed with NaN omitted):
   • Spearman $\rho$ (global): Spearman rank correlation $\rho$ evaluated on the entire dataset.
@@
   • Top-64/top-$10\%$: number of top-64 models ranked by zero-cost scoring metric within top-$5\%$ performing models.
-2.1.1 NAS-Bench-201
-The results for overall epsilon performance on NAS-Bench-201 are given in Table 1 along with other zero-cost NAS metrics.
-The Kendall $\tau$ score is not reported in Abdelfattah et al.
-(2021), but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation.
-We use the data provided by Abdelfattah et al.
-(2021) to evaluate their Kendal $\tau$.
-Note that our results differ from the original paper for some evaluation scores.
+5 Results
+5.1 NAS-Bench-201
+The results for overall epsinas performance on NAS-Bench-201 are given in Table 2 along with other zero-cost NAS metrics.
+The Kendall $\tau$ score is not reported in [1], but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation.
+We use the data provided by [1] to evaluate their Kendal $\tau$.
+Note that our results differ from the original paper in terms of evaluation scores.
 In such cases, we indicate the original values between brackets.
 In particular, there is a discrepancy in computing the values in the last column, Top-64/top-$5\%$, while the rest of the results are consistent.
-Figure 6 in Appendix suggests that our calculations are correct.
+Figure 8 in the Appendix suggests that our calculations are correct.
 Table 2: Zero-cost metrics performance for the NAS-Bench-201 search space with its three datasets: CIFAR-10, CIFAR-100 and ImageNet16-120.
-We give the original values from Abdelfattah et al.
-(2021) for reference between brackets.
-We highlight the best-performing metrics in bold.
+We give the original values from [1] for reference between brackets.
+We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
 Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 fisher 0.36 -0.38 0.26 -0.24 4.99 ( 5) 0 (0)
-epsilon 0.87 0.55 0.70 0.40 67.39 59
+nparams 0.75 0.57
+zico 0.80 0.61
+$\xi$-gsnr 0.845 0.661
+epsinas 0.87 0.50 0.69 0.36 67.06 62
 CIFAR-100
@@
 fisher 0.39 -0.15 (-0.16) 0.28 -0.10 4.22 ( 4) 0 (0)
-epsilon 0.90 0.59 0.72 0.43 81.24 62
+nparams 0.73 0.55
+zico 0.81 0.61
+$\xi$-gsnr 0.840 0.658
+epsinas 0.89 0.63 0.72 0.45 81.33 62
 ImageNet16-120
@@
 fisher 0.33 0.02 0.25 0.01 4.61 ( 5) 0 (0)
-epsilon 0.85 0.53 0.67 0.37 71.51 59
+nparams 0.69 0.52
+zico 0.79 0.60
+$\xi$-gsnr 0.793 0.608
+epsinas 0.85 0.53 0.67 0.37 71.51 59
 For NAS-Bench-201, we also report average performance when selecting one architecture from a pool of $N$ random architectures.
 The statistics are reported over $500$ runs.
-Table 2 compares epsilon to other trainless metrics.
-Note that te-nas starts with a super-network composed of all the edges and operators available within the space.
-In this case, $N$ is not applicable, and the performance can not be improved.
-In principle, other methods’ performance should improve with higher $N$ values.
-Table 2: Comparison of the trainless metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets.
-On the top, we list the best-performing methods that require training: REA (Real et al.
-(2019)), random search, REINFORCE (Williams (1992)), BOHB (Falkner et al.
-(2018)), DARTS- (Chu et al.
-(2020)).
+Table 3 compares epsinas to other zero-cost metrics.
+Table 3: Comparison of the zero-cost metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets.
+On the top, we list the best-performing methods that require training: REA [38], random search, REINFORCE [49], BOHB [13], DARTS- [7].
 We report the average best-achieved test accuracy over $500$ runs, with $1{,}000$ architectures ($100$ for grad_sign) sampled from the search space at random.
-For tenas, the results are reported for $4$ random seeds.
 Random and optimal performances are given as a baseline.
@@
 Random n/a $84.11\pm 11.71$ $87.40\pm 11.94$ $61.57\pm 11.35$ $61.67\pm 11.35$ $33.97\pm 8.68$ $33.67\pm 8.98$
-Trainless (N=1000)
+Zero-cost (N=1000)
 naswot $3.3$ $89.69\pm 0.73$ $92.96\pm 0.81$ $69.86\pm 1.21$ $69.98\pm 1.22$ $43.95\pm 2.05$ $44.44\pm 2.10$
@@
 fisher n/a $87.01\pm 2.31$ $87.36\pm 2.23$ $63.54\pm 5.69$ $63.67\pm 5.62$ $26.70\pm 10.83$ $29.56\pm 10.83$
-epsilon $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
-Trainless (N=100)
+epsinas $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
+Zero-cost (N=100)
 cv $0.05$ $84.89\pm 6.39$ $91.90\pm 2.27$ $63.99\pm 5.61$ $64.08\pm 5.63$ $38.68\pm 6.34$ $38.76\pm 6.62$
 grad_sign n/a $89.84\pm 0.61$ $93.31\pm 0.47$ $70.22\pm 1.32$ $70.33\pm 1.28$ $42.07\pm 2.78$ $42.42\pm 2.81$
-epsilon $0.003$ $\boldsymbol{90.44}\pm\boldsymbol{0.97}$ $\boldsymbol{93.39}\pm\boldsymbol{0.82}$ $\boldsymbol{70.85}\pm\boldsymbol{1.30}$ $\boldsymbol{71.00}\pm\boldsymbol{1.26}$ $\boldsymbol{44.03}\pm\boldsymbol{2.02}$ $\boldsymbol{44.20}\pm\boldsymbol{2.04}$
-Trainless (all space)
-tenas $0.5$ $93.9\pm 0.47$ $71.24\pm 0.56$ $42.38\pm 0.46$
-Comparing the results for epsilon with other zero-cost NAS metrics, we can see that it outperforms them by a good margin.
-Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar figures for other methods can be found in Appendix, Figure 6).
-However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsilon development.
+epsinas $0.003$ $90.44\pm 0.97$ $93.39\pm 0.82$ $70.85\pm 1.30$ $71.00\pm 1.26$ $44.03\pm 2.02$ $44.20\pm 2.04$
+Comparing the results for epsinas with other zero-cost NAS metrics, we can see that it shows surprisingly good performance given its conceptual simplicity.
+Importantly, it is consistently better than the naive estimator nparams.
+Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar plots for other methods can be found in the Ablation section, Figure 8).
+However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsinas development.
 Refer to captionRefer to captionRefer to caption
-Figure 1: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets.
+Figure 1: Zero-cost NAS epsinas metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets.
 The horizontal axis shows test accuracy upon training.
@@
 The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).
-2.1.2 NAS-Bench-101
-We use NAS-Bench-101 space to confirm that the success of epsilon metric in the previous section is not due to overfitting the NAS-Bench-201 search space and to see how it applies to a vaster search space.
-Table 3 together with Figure 2 confirm that it performs reasonably well on NAS-Bench-101, too.
+5.2 NAS-Bench-101
+We use the NAS-Bench-101 search space to verify that the effectiveness of the epsinas metric, as demonstrated in the previous section, is not merely a result of overfitting to the NAS-Bench-201 space.
+It also allows us to assess its performance in a broader search space.
+As shown in Table 4 and Figure 2, epsinas performs well on NAS-Bench-101, once again surpassing the nparams proxy.
 Table 4: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset.
-Values from Abdelfattah et al.
-(2021) are given for reference between brackets.
-We highlight the best-performing metrics in bold.
+Values from [1] are given for reference between brackets.
+We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
 Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 fisher -0.26 -0.07 (0.07) -0.18 -0.05 2.65 (3) 0 (0)
-epsilon 0.62 0.12 0.44 0.08 40.33 10
+nparams 0.43 0.31
+zico 0.63 0.46
+$\xi$-gsnr 0.615 0.434
+epsinas 0.62 0.12 0.44 0.08 40.33 10
 Refer to captionRefer to caption
-Figure 2: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset.
+Figure 2: Zero-cost NAS epsinas metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset.
 The horizontal axis shows test accuracy upon training.
@@
 The figure shows $423{,}624$ and $14{,}322$ networks for NAS-Bench-101 and NAS-Bench-NLP, respectively (excluding architectures with NaN scores).
-2.1.3 NAS-Bench-NLP
+5.3 NAS-Bench-NLP
 Both NAS-Bench-201 and NAS-Bench-101 are created to facilitate NAS in image recognition.
 They operate convolutional networks of very similar constitutions.
-To truly probe the generalisability of the epsilon metric, we test it on NAS-Bench-NLP.
+To truly probe the generalisability of the epsinas metric, we test it on NAS-Bench-NLP.
 Both input data format and architecture type differ from the first two search spaces.
-Unfortunately, Abdelfattah et al.
-(2021) provides no data for NAS-Bench-NLP, disabling us from using their results for calculations.
-Therefore, in Table 4, we give only values provided in the paper together with our epsilon metric (data for ficher is absent).
-We want to note that, unlike accuracy, perplexity used for language-related ML problems should be minimised.
-Therefore, the signs of correlations with scoring metrics should be reversed, which is not the case for numbers given in Abdelfattah et al.
-(2021).
-Table 4: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We highlight the best-performing metrics in bold.
+Unfortunately, [1] provides no data for NAS-Bench-NLP, disabling us from using their results for calculations.
+Therefore, in Table 5, we give only values provided in the paper together with our epsinas metric (data for ficher is absent).
+It’s important to note that, unlike accuracy, perplexity – commonly used in language-related machine learning problems – should be minimized.
+Thus, the signs of correlations with scoring metrics must be reversed, which differs from the results reported in [1].
+Table 5: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset.
+We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
 Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 snip -0.19 -0.02 10
-epsilon -0.34 -0.12 -0.23 -0.08 24.87 11
+epsinas -0.34 -0.12 -0.24 -0.08 26.23 13
+5.4 Computational cost
+One of the primary purposes of zero-shot proxies is to reduce the computational cost of the NAS process.
+Since epsinas only requires two forward passes through the data and does not involve gradient computation, it is efficient regarding both time and memory usage.
+Figure 3 presents the results of tests conducted using NAS-Bench-101.
+As anticipated, memory usage and processing time scale linearly with the number of parameters and FLOPs.
+It also demonstrates that epsinas can be computed using a CPU within 30 seconds for the largest NAS-Bench-101 networks consisting of several million parameters.
+Figure 3: Computational efficiency of the zero-cost NAS epsinas metric for the NAS-Bench-101 search space.
+The results are based on the evaluation of $10{,}000$ networks from the NAS-Bench-101 benchmark, divided into $20$ bins according to the number of parameters.
+The shaded area represents one standard deviation within each group.
+FLOPs are reported for a batch size of $1$, while memory usage and processing time are evaluated with a batch size of $256$.
+Comparing epsinas to other zero-cost metrics is not straightforward, as different methods use different search spaces, computational setups, and batch sizes.
+For instance, zico reports evaluation times using MobileNetV2 [41] with a budget of $1,000$ MFLOPs and a batch size of $128$.
+The authors noted that evaluating $10^{5}$ architectures required $10$ GPU hours.
+MobileNetV2 networks range from $1.7$M to $6.9$M parameters, larger than NAS-Bench-201 but smaller than NAS-Bench-101.
+To provide a closer comparison, we evaluated epsinas on NAS-Bench-101 with a batch size of $128$, limiting architectures to $6.9$M parameters.
+Under these conditions, evaluating $10^{5}$ architectures took approximately $7$ GPU hours, with the epsinas computation requiring only $2$ GPU hours; the remaining time was spent on model loading.
+While this is not a direct comparison, epsinas does not involve gradient computation, so it is reasonable to expect it to be faster per architecture evaluation. The results support this expectation.
+5.5 Correlation with nparams
+As noted in previous studies [35, 25, 48], one of the best naive evaluators of a network’s performance is the number of parameters, nparams.
+Generally, the more parameters a network has, the better we expect its predictive performance to be.
+However, bigger does not always mean better: large networks are susceptible to overfitting and can be prohibitively expensive to train [8, 22].
+Moreover, research on network pruning has shown that, with the right architecture, smaller subnetworks can achieve comparable or even superior performance to their larger counterparts [15].
+Hence, when seeking an optimal evaluation proxy, it is essential to ensure that it can effectively identify smaller networks.
+Figure 4 illustrates the relationship between epsinas and the number of parameters, showing no strong preference for larger network sizes.
+Figure 4: Visualisation of the relationship between epsinas score and the number of parameters in a network, nparams for CIFAR-10 data of NAS-Bench-201 benchmark.
 5.6 Integration with other NAS methods
-While it is possible to utilise zero-cost metrics independently, they are often implemented within other NAS.
-Here we provide examples of random search and ageing evolution algorithms when used in tandem with epsilon.
-Similarly to Abdelfattah et al.
-(2021), we compare random search performance with and without warm-up.
-First, we create a warm-up pool of $3{,}000$ architectures.
-Then, during the first $64$ steps of random search, the algorithm picks networks from the pool based on the highest trainless score and accordingly updates the best test performance.
-For ageing evolution, the same principle applies.
-En plus, we report the results of implementation where every next parent is decided based on the highest epsilon score (”move”).
-In other words, the trainless scoring metric replaces validation accuracy in move mode.
-Finally, the child is created by parent mutation (within an edit distance of $1$) and added to the pool.
-Finally, the oldest network is removed from the pool.
-For both described algorithms, we run the procedure until the number of trained architectures reaches $300$ and perform $100$ random rounds.
-Figure 3 shows that epsilon metric leads to considerable improvements in terms of time and precision.
-The best performance is achieved in combination with a warm-up.
-Figure 5 assembles warm-up performances for several trainless metrics.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 3: epsilon integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
+While zero-cost metrics are robust on their own, they benefit from being integrated with other NAS algorithms.
+[32] showed that combining a zero-cost metric based on Jacobian covariance, similar to [34], with REA [38] speeds up evolutionary search and helps find top-performing architectures.
+This work demonstrates how the epsinas metric can improve random search and ageing evolution algorithms.
+This work follows the implementation of [1].
+Random search Each new architecture is selected randomly and evaluated based on test accuracy. The final test performance is reported without using any auxiliary metrics.
+Random search with warm-up A warm-up pool of $3{,}000$ randomly sampled architectures is first ranked using epsinas.
+Next, models are ”trained” in order of decreasing epsinas value by retrieving their test accuracies.
+Training stops once the budget of $300$ architectures is reached.
+Ageing evolution We implement the basic ageing evolution algorithm from [38].
+The process starts with a randomly sampled pool of $64$ architectures.
+Parents are selected based on test accuracy, and a child is generated through mutation (with an edit distance of $1$).
+The child is added to the pool, while the oldest architecture is removed.
+This process repeats until $300$ architectures have been trained.
+Ageing evolution with warm-up A warm-up phase begins with $3{,}000$ sampled architectures, which are ranked using the epsinas metric.
+The $64$ architectures with the highest epsinas values form the initial pool.
+The standard ageing evolution process continues, where parents are selected based on test accuracy.
+Ageing evolution with move Instead of using test accuracy for parent selection, architectures are assessed based on their epsinas scores.
+We conduct $100$ random runs for each algorithm and report the mean and standard deviation.
+Figure 5 demonstrates that the epsinas metric significantly improves both efficiency and accuracy, with the best results achieved when combined with a warm-up.
+Figure 7 compares warm-up performances across various zero-cost metrics.
+Figure 5: epsinas integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
 6 Ablation studies
-While our metric is relatively straightforward to implement, it does require several hyperparameters to be set. In this section, we provide the ablations studies’ results and analyse them.
+While our metric is straightforward to implement, it depends on several hyperparameters. In this section, we present and analyse the results of our ablation studies.
 6.1 Weights
-To compute the epsilon metric, we initialise the networks with two distinct constant shared weights. The question is: how to fix their values, and how does this choice affect the whole method?
+To compute the epsinas metric, we initialise the networks with two distinct constant shared weights. The question is: how do we fix their values, and how does this choice affect the whole method?
 To answer this question, we ran a series of tests with multiple pairs of weights.
-In pairs, it makes sense to set the first weight always greater than the second one.
-This consideration stems from the fact that epsilon is based on the mean absolute difference between two initialisations.
+It makes sense to set the first weight always greater than the second one in pairs.
+This consideration stems from the fact that epsinas is based on the mean absolute difference between two initialisations.
 Therefore, it will output the same score for [$w_{1}$, $w_{2}$] and [$w_{2}$, $w_{1}$] combinations and will be exactly zero in the case where the two weights are equal.
-Tables 5, 6, 7 summarises the results of our tests.
+Tables 6, 7, 8 summarise the results of our tests.
 For the NAS-Bench-201 search space, the outcomes are very similar between the datasets, and the minor differences do not affect the conclusions drawn from this data.
-Table 5: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset.
+Table 6: Zero-cost metrics performance was evaluated using the NAS-Bench-101 search space and the CIFAR-10 dataset.
 We start with $5{,}000$ randomly picked architectures.
-The number of architectures with non-NaN values are given in the second column.
+The number of architectures with non-NaN values is given in the second column.
 Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 0.1, 1 5000 -0.47 0.04 -0.32 0.03 0.20 0
-Table 6: Zero-cost metrics performance evaluated on NAS-Bench-201 search space, CIFAR-10 dataset.
+Table 7: Zero-cost metrics performance was evaluated in the NAS-Bench-201 search space and the CIFAR-10 dataset.
 We start with $5{,}000$ randomly picked architectures.
-The number of architectures with non-NaN values are given in the second column.
+The number of architectures with non-NaN values is given in the second column.
 Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 0.1, 1 4907 0.01 0.26 -0.03 0.16 0.84 0
-Table 7: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset.
+Table 8: Zero-cost metrics performance was evaluated using the NAS-Bench-NLP search space and the PTB dataset.
 We start with $5{,}000$ randomly picked architectures.
-The number of architectures with non-NaN values are given in the second column.
+The number of architectures with non-NaN values is given in the second column.
 Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
 0.1, 1 652 -0.27 0.20 -0.19 0.13 10.61 7
-The results show that there is indeed some variation between different initialisations.
-Two effects impact performance.
-Firstly, the more significant the difference between the two weights, the higher the correlation between epsilon and trained performance.
-It makes sense since if the weights are too close to each other, the difference between the outputs becomes minimal, and the distinction between architectures in terms of epsilon is no longer possible.
+The results indicate that variations in initialisation do affect performance.
+Two key factors influence this impact.
+First, a more significant difference between the two weights leads to a higher correlation between epsinas and trained performance.
+It is intuitive – when the weights are too similar, the output difference becomes minimal, making it difficult to distinguish between architectures based on epsinas.
 Secondly, we can see that too low and too high weights result in fewer architectures with non-NaN scores.
-Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsilon metric.
+Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsinas metric.
 Naturally, the deeper the network is, the more signal attenuation is, and the higher the probability of NaN score.
-We remind that prior to epsilon computation, the outputs are normalised to the same range $[0,1]$, which means the value of the metric per se does not become smaller for deeper networks.
+We remember that before epsinas computation, the outputs are normalised to the same range $[0,1]$, which means the metric value per se does not become smaller for deeper networks.
 This extinction effect is due to the limited sensitivity of float type.
 Theoretically, with infinite float sensitivity, there would be no extinction and hence no effects on the choice of the weights.
-Table 8: Optimal weights for epsilon evaluation for three search spaces and their datasets.
+Table 9: Optimal weights for epsinas evaluation for three search spaces and their datasets.
 Search space Dataset Optimal weights
@@
 NAS-Bench-NLP PTB $[10^{-5},10^{-3}]$
-As a rule of thumb, for every new ML problem, we suggest running epsilon evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs.
+As a rule of thumb, for every new ML problem, we suggest running epsinas evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs.
 While we acknowledge that this procedure is subjective, importantly, it does not require any training.
@@
 The choice to initialise the weights with constant values provides a ground for discussion.
-On the one hand, constant shared weights guarantee the same initialisation regardless of the structure; the noise from the randomness of other initialisations is not obstructing the effect of the neural geometry.
-It would be possible to reduce the amount of noise by increasing the number of independent initialisations with distinct random seeds similar to Gracheva (2021), but this increases the algorithm run-time proportionally.
-Table 9 gives the results of the ablation studies when we compute epsilon based on two initialisations.
-Table 9: Initialisations ablation studies. The metric is computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
-Initialisation Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
+On one hand, using constant shared weights ensures a consistent initialisation across different structures, eliminating the noise introduced by random initialisations.
+It allows us to isolate and analyse the impact of neural geometry more effectively.
+Averaging results over multiple runs with different random seeds, as in [18], can reduce the randomness noise but increases computational time proportionally.
+Table 10 presents the results of our ablation studies, where epsinas is computed using different initialisation schemes.
+The correlation coefficients are based on the epsinas scores from two initialisations but using standard initialisation methods instead of constant shared weights.
+For every initialisation scheme, we only modify the convolutional layers.
+The rest of the architecture is initialised as per the original benchmark space: constant initialisation with ones for the batch normalisation, normal uniform for the linear layer, and biases set to zero.
+Table 10: Initialisations ablation studies. The metric is computed over the first $5{,}000$ architectures in the NAS-Bench-201 benchmark.
+Initialisation Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
 global top-10% global top-10% top-10% top-5%
-Uniform 15284 0.14 -0.17 0.09 -0.11 3.07 7
-Normal 15284 0.44 -0.29 0.30 -0.20 20.07 4
-Kaiming uniform 15284 0.32 0.11 0.21 0.07 27.40 26
-Kaiming normal 15284 0.49 -0.29 0.33 -0.20 17.86 7
-Orthogonal 15284 0.41 -0.20 0.28 -0.13 15.11 1
+Uniform -0.02 -0.01 -0.01 -0.00 7.13 13
+Normal 0.08 -0.20 0.05 -0.13 15.83 4
+Kaiming uniform 0.07 -0.17 0.05 0.12 13.60 2
+Kaiming normal 0.07 -0.13 0.05 -0.09 12.49 1
+Orthogonal 0.07 -0.16 0.05 -0.10 11.71 6
 6.3 Batch size
@@
 The tests are done on the NAS-Bench-201 search space with its three datasets.
-For each batch, we report medians with $25$ and $75$ percentiles over $10$ runs.
-We evaluate epsilon on $5{,}000$ architectures.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Refer to caption Refer to caption
+We report medians with $25$ and $75$ percentiles over $10$ runs for each batch.
+We evaluate epsinas on $5{,}000$ architectures.
 Figure 6: Batch size ablation study for NAS-Bench-201 search space, CIFAR-10 (left), CIFAR-100 (centre) and ImageNet16-120 (right) datasets.
-Figure 4 show that, as expected, the larger the batch size, the better and more stable the performance of epsilon metric.
+Figure 6 shows that, as expected, the larger the batch size, the better and more stable the performance of epsinas metric.
 There is almost no improvement for batches over $256$.
@@
 6.4 Embedding initialisation
-In order to verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
+To verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
   • uniform positive 0.1: random uniform from ranges [0, 0.1]
@@
 Care should be taken when initialising the networks containing the embedding layer: if embedding is initialised with all constants, there is no difference between the embedded input.
-The performance of our metric, in this case, will be analogous to that with a batch size of one.
-Our ablation studies show that as soon as the embedding is initialised with non-constant weights, it does not significantly influence the outcomes.
+In this case, our metric’s performance will be analogous to that with a batch size of one.
+Our ablation studies show that it does not significantly influence the outcomes as soon as the embedding is initialised with non-constant weights.
 6.5 Synthetic data
@@
   • grey scale images: images within the batch are solid colour ranging from black to white
-  • random normal: images are filled with random values following normal distribution with $[\mu,\sigma]=[0,1]$
+  • random normal: images are filled with random values following a normal distribution with $[\mu,\sigma]=[0,1]$
   • random uniform: images are filled with random values following uniform distribution with $[\mu,\sigma]=[-1,1]$
@@
 Curiously, greyscale images, filled with constant values, show the closest results to the actual data.
-Note that epsilon metric with greyscale data outperforms synflow.
+Note that epsinas metric with greyscale data outperforms synflow.
 It is an essential achievement because synflow does not use input data and is, therefore, data independent.
-Our results show that epsilon has the potential to be used with no data whatsoever.
+Our results show that epsinas has the potential to be used with no data whatsoever.
 Table 12: Synthetic data tests with CIFAR-10 dataset, NAS-Bench-201 search space. Tests are based on the evaluation of $5{,}000$ architectures.
@@
 Random uniform (+) 0.61 0.17 0.43 0.16 16.22 $4$
-4 Visualisation of other zero-cost NAS metrics
-4.1 When integrated with other NAS methods
-In Section 2.2, we show how epsilon metric improves the performance of ageing evolution and random search when used for warming up.
-Figure 5 compares epsilon integration to other metrics from Abdelfattah et al.
-(2021).
-For both EA ans RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds.
+7 Discussion
+Our experiments show that despite its simplicity, epsinas metric shows surprisingly good performance across different benchmark datasets.
+It shows the most consistent performance with NAS-Bench-201, suggesting that this benchmark defines a relatively easy-to-navigate search space.
+Tables 2, 4, 5 demonstrates that epsinas outperforms most of other zero-cost metrics in terms of Spearman and Kendal correlations.
+The differences are statistically significant, with p-values between epsinas and other metrics below 0.001.
+Notably, we highlight the key differences between our previous work (cv metric, [18]) and the current approach.
+The two methods differ in how they de-emphasize weights and the type of outputs used for statistical computation.
+Using two fixed weight initialisations more effectively isolates the influence of network geometry compared to averaging over a hundred random initialisations.
+Additionally, the epsinas metric is significantly more efficient, reducing computation time by a factor of fifty.
+We also compare the performance of epsinas with its conceptual counterparts, zico and $\xi$-gsnr.
+While the metrics are based on similar principles, zico and $\xi$-gsnr use gradients instead of raw outputs, which appears to limit their performance.
+This difference may also be attributed to random initialisation, as opposed to the shared constant weight initialisation used by epsinas.
+In all cases, epsinas consistently outperforms the naive nparams proxy, which relies solely on the total number of parameters in a given network.
+It is important to note that we used NAS-Bench-201 to develop our metric, so the findings from this part of the study may be biased.
+To test the generalizability of the metric, we applied it to the NAS-Bench-101 benchmark.
+As shown in Table 4, epsinas continues to demonstrate strong performance.
+The peculiar vertical lines in the results are due to clustering based on test accuracy.
+Since each architecture in NAS-Bench-101 is trained with three random seeds, the four clusters represent cases where training diverges for certain random initialisations.
+Using more random seeds would likely reveal a smoother pattern and stronger correlations.
+By design, we expect our metric to be independent of the specific neural architecture structure. To confirm this claim, we show its applicability to architectures to RNN models with NAS-Bench-NLP.
+While the statistics in Table 5 may not seem extraordinary at first glance, Figure 2 demonstrates a visible trend of improved architecture selection as the epsinas metric value increases.
+This positive trend is particularly noteworthy, especially considering the noise inherent to the benchmark.
+Factors such as the limited sample size of networks within a vast architecture space, variations in training hyperparameters, dropout rates, and other variables can contribute to performance inconsistencies in NAS.
+Despite these challenges, the trend is strong and consistent enough to confidently conclude that the epsinas metric shows great potential for effectively evaluating recurrent-type architectures.
+While epsinas metric shows solid empirical performance, the underlying reasons for this remain unclear.
+There are several hints towards its understanding.
+First, mathematically, epsinas represents the difference in the output distribution shapes between initialisations.
+The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry.
+By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
+Second, according to the weight ablation studies (Section 6), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive output explosion or vanishing.
+Therefore, epsinas measures the amplitude of the outputs’ distribution shape change due to geometry.
+Finally, we see that grey-scale solid images work reasonably well as inputs during the synthetic data studies.
+The distribution over the input samples is uniform, which makes it easier to track the changes in the distribution shape as the signal propagates through the network.
+7.1 Future directions
+While the presented epsinas metric demonstrates solid performance, several open questions remain.
+One of the key directions for future development is to assess its applicability to more types of neural topologies, such as transformer [3, 54], generative adversarial [17] or graph neural networks [39], and to test it with hierarchical search spaces [6, 39, 29, 3].
+It is also interesting to evaluate the metric in combination with other advanced NAS algorithms, such as multi-objective optimization [53], differentiable architecture search [50], Bayesian optimization [43, 42], or within the Parametric Zero-Cost Proxies [10] framework.
+The metric’s performance is influenced by various hyperparameters, including weight values, batch size, and initialization scheme.
+While we conducted ablation studies on their individual effects, potential interactions between these hyperparameters remain unexplored, making it important to investigate their joint impact.
+Additionally, the current weight setup is based on empirical choices and would benefit from a solid theoretical foundation.
+A deeper theoretical study of the metric is particularly needed given its similarities to the zico metric.
+For instance, it remains unclear why the mean absolute error over raw outputs outperforms the deviation over gradients and how this behaviour relates to the initialization scheme.
+Furthermore, examining the performance of the zico metric when initialized with constant shared weights could provide further insights.
+8 Conclusion
+This work presents a simple zero-cost NAS scoring metric, epsinas, which measures how much a network’s outputs change when initialized with two different constant shared weights.
+We compute it by performing two forward passes and calculating the normalized MAE between the raw outputs.
+For a simple metric, epsinas shows an unexpectedly high correlation with network performance in CNNs and RNNs, outperforming most existing zero-shot NAS proxies.
+We evaluate the epsinas metric on three widely-used NAS search spaces: NAS-Bench-201, NAS-Bench-101, and NAS-Bench-NLP.
+Our results show strong performance across these search spaces.
+The metric is computationally efficient, requiring only a fraction of a GPU second to evaluate a single architecture, with minimal memory demand, as it does not require gradient computation or labelled data.
+epsinas can be run on a CPU, making it broadly accessible.
+While small to medium-sized networks can be processed efficiently on a CPU, larger architectures may require GPU acceleration to handle forward passes more quickly, especially when evaluating multiple candidate architectures in succession.
+Optimizing CPU-GPU workload distribution could help balance energy consumption and speed in resource-constrained environments such as mobile and embedded systems.
+In our implementation, we have already incorporated a degree of heterogeneous computing by offloading output data from the GPU to the CPU for statistical computation.
+Such optimisations could be valuable in domains where rapid evaluation of complex architectures is necessary.
+Experimental results demonstrate that epsinas applies to convolutional networks and generalizes well to recurrent networks—with careful consideration for embedding layer initialization.
+The method can also be readily incorporated into other NAS frameworks, such as random search and evolutionary algorithms.
+The main downside of our metric is that the optimal weight values need to be set separately for different search spaces and neural topologies.
+We found that using both high and low weight values helps highlight differences in initialization, leading to better correlations with final performance.
+While the weights should cover both extremes to test the architecture’s robustness, it’s unclear how to automate this selection process.
+These results offer an intriguing clue about the inner workings of neural networks.
+We hope that epsinas ’s performance can be improved with a better theoretical understanding of the optimal weight setup.
+We believe that further development of this method will lead to a faster and computationally efficient NAS and a clearer understanding of network prediction and generalization mechanisms.
+These results might offer some interesting clues about how neural networks work.
+We are hopeful that with a better theoretical understanding of the optimal weight setup, epsinas ’s performance can improve, and we believe that further development of this method could help us better understand network prediction and generalization.
+Declarations
+Competing interests
+The authors have no competing interests to declare that are relevant to the content of this article.
+Funding
+No funding was received for conducting this study.
+Code availability
+The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
+Appendix A Visualisation of other zero-cost NAS metrics
+A.1 When integrated with other NAS methods
+In Section 5.6, we show how epsinas metric improves the performance of ageing evolution and random search when used for warming up.
+Figure 7 compares epsinas integration to other metrics from [1].
+For both EA and RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds.
 The warm-up pool contains $3000$ randomly selected architectures.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Refer to caption Refer to caption
 Figure 7: Various zero-cost NAS metrics’ performance when used as a warm-up for ageing evolution (top) and random search (bottom) for three datasets from the NAS-Bench-201 search space.
@@
 The warm-up population is $3000$ architectures.
-4.2 Correlation with accuracy
-In Abdelfattah et al.
-(2021) the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding.
-Here we provide visualisations for two search spaces built on the data provided by the authors.
+A.2 Correlation with accuracy
+In [1], the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding.
+Here, we provide visualisations for two search spaces built on the data provided by the authors.
 Refer to captionRefer to captionRefer to caption
@@
 Refer to captionRefer to captionRefer to caption
-Figure 6: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from Abdelfattah et al.
-(2021).
+Figure 8: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from [1].
 Each dot corresponds to an architecture; the darker the colour, the more parameters it contains.
-Figure represents search space of $15{,}625$ networks (excluding architectures with NaN scores).
-5 Discussion
-The epsilon metric performs surprisingly well on various benchmark datasets.
-Similarly to other NAS algorithms, it shows the most robust performance with the NAS-Bench-201 benchmark.
-This fact indicates that the set of architectures comprising NAS-Bench-201 forms a relatively easy-to-navigate landscape regardless of the metric.
-Still, epsilon shows the best performance among trainless metrics.
-Table 1 shows that synflow shows comparable performance in terms of correlations on the ImageNet-16-120 dataset.
-However, statistically epsilon selects, on average, a better-performing architecture, which can also be confirmed visually by comparing Figures 1 and 6.
-Notably, we would like to highlight the difference with our previous work (cv metric, Gracheva (2021)).
-The two methods differ in how the weights are de-emphasised and what outputs we use for statistics computation.
-Two constant shared weight initialisations better bring out the effect of the bare geometry compared to a hundred random initialisations.
-Moreover, epsilon metric takes significantly less time to compute, being fifty-fold faster.
-It is important to note that we used NAS-Bench-201 for the metric development; therefore, the results of this part of the study might be biased.
-To confirm the generalisability of the metric, we applied it to the NAS-Bench-101 benchmark.
-Again, Table 3 shows strong performance.
-The peculiar vertical lines are due to clustering in terms of test accuracy.
-Since every architecture in NAS-Bench-101 is trained with $3$ random seeds, the four clusters correspond to situations when the training diverges for some random initialisations.
-More random seeds would reveal a smoother picture and better correlations.
-Again, we observe comparable performance between synflow and epsilon scoring metrics.
-We could only confirm some of the results from Abdelfattah et al.
-(2021) and suppose that some statistics might have been miscalculated.
-By design, we expect our metric to be independent of the specific neural architecture structure.
-To confirm this claim, this is essential to show its applicability to architectures other than CNNs.
-Our work tests epsilon applicability to RNN models with NAS-Bench-NLP.
-While statistics in Table 4 are not exceptional, Figure 2 shows a visible trend towards better architectures with increasing metric value.
-Significant noise levels might stem from the characteristics of the benchmark itself: factors like a relatively small sample of networks given a vast space, choice of training hyperparameters, dropout rates, and others may distort NAS performance.
-Nonetheless, the trend is strong enough to conclude that epsilon metric can apply to recurrent type architectures.
-While epsilon metric shows solid empirical performance, the underlying reasons for this are unclear.
-There are several hints towards its understanding.
-First, mathematically, epsilon represents the difference in the output distribution shapes between initialisations.
-The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry.
-By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
-Second, according to the weight ablation studies (Section 3), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive outputs explosion or vanishing.
-Therefore, epsilon measures the amplitude of the outputs’ distribution shape change due to geometry.
-Finally, we see that grey-scale solid images work reasonably well as inputs during the synthetic data studies.
-The distribution over the input samples is uniform, which makes it easier to track the changes in the distribution shape as the signal propagates through the network.
-6 Conclusion
-This work presents a novel zero-cost NAS scoring metric epsilon.
-It allows evaluating a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations.
-The value of epsilon reflects how neural network outputs’ distribution changes between low and high constant shared weight initialisations.
-It shows that the higher this difference, the better the network will perform upon training.
-It takes a fraction of a second to evaluate a single architecture and does not require gradients computation (low demand on memory) or labels, and can run on a CPU.
-We provide experimental evidence that the introduced epsilon metric generalises to various architecture types, such as convolutional or recurrent neural networks (care should be taken with embedding layer initialisation).
-This work evaluates epsilon on three staple NAS search spaces: NAS-Bench-201, NAS-Bench-101 and NAS-Bench-NLP.
-It sets the state-of-the-art among zero-cost NAS metrics and can be easily implemented within other methods, such as random search or evolutionary algorithms.
-The only significant disadvantage of the method is that it requires the choice of constant weight values used for initialisation.
-Our tests show that it must be set up individually for each search space.
-We plan to automate the weight selection process in our future work.
-Overall, while a coherent theoretical foundation of epsilon is missing, we consider the empirical results compelling.
+The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).